OpenStudy (anonymous):

Are there 2 solutions for the system shown below? x^2+y^2=25 x-y^2=-5

OpenStudy (raden):

2nd equation : x-y^2=-5 x = y^2 - 5 subtitute the equation above into 1st equation : (y^2 - 5)^2+y^2=25 y^4 -9y^2 = 0 y^2(y^2 - 9) = 0 y^2 (y+3)(y-3) = 0 now how many solution make y be equals zero ?

OpenStudy (anonymous):

just one equation makes y equal to zero, right?

OpenStudy (radar):

Be sure and test possible answers. Some of these may not work for the system: (4, -3) (4, 3) (-5,0) (5,0)

OpenStudy (raden):

setting each factors bellow to zero, then solve for y y^2 (y+3)(y-3) = 0 y^2 = 0 solve for y ? y + 3 = 0 solve for y ? y - 3 = 0 solve for y ? how many different of y's ?

OpenStudy (anonymous):

only y^2=0 gives me 0. y+3 gives me -3 and y-3 gives me positive 3

OpenStudy (raden):

btw, (5,0) cant be satisfies for the 2nd equation :). @radar x-y^2=-5 5 - 0^2 = -5 5 = -5 (incorrect)

OpenStudy (raden):

yes, you are right, @ali1029

OpenStudy (radar):

That was my point,.....

OpenStudy (anonymous):

so then there is only 1 solution possible?

OpenStudy (raden):

actually, if you have 3 y's as solutions then you have 3 x's too. therefore, we have 3 solutions for (x,y)

OpenStudy (anonymous):

oh I see, that completely makes sense! thank you :)

OpenStudy (raden):

welcome